A gas contain in a flask of 500ml capacity was allowed to expand into another flask of volume 600ml connecting through a valve and at the same temperature,if the final pressure of the gas is 1.01×10^5pa,When the valve is open,what was the pressure of the gas in d 500mmgh?The volume of the flask is negligible


1
Expert's answer
2021-10-18T11:02:44-0400

"V_1 = 500\text{ mL} = 500 × 10^{-6}\text{ m³}\\
V_2 = 600\text{ mL} = 600 × 10^{-6}\text{ m³}\\
V_t = (500 + 600) \text{ mL}= 1100\text{ mL} = 1100 × 10^{-6}\text{ m³}\\
P_1 =\ ?\\
P_2 = P_t= 1.01 × 10⁵\text{ Pa}"

volume of valve is negligible

"P_1V_1 = P_tV_t\\
P_1 × 500 = 1.01 × 10⁵ × 1100\\
P_1 × 500 = 1111 × 10^5\\
P_1 = 2.22 × 10⁵ \text{ Pa}"